OpenStudy (anonymous):

Henry Devine bought a new dishwasher for $320. He paid $20 down and made 10 monthly payments of $34. What actual yearly interest rate did Henry pay? A) 29.09 B) 68.75 C) 34.38 D) 14.55 I need help figuring the problem out.

OpenStudy (anonymous):

all you have to do is 34*10 and your answer would be c\[34*10\]

OpenStudy (mathstudent55):

You need to find how much total interest he paid. Then how much money he financed. Also, you need to know how many months from the purchase date it took him to pay.

OpenStudy (catlover5925):

Henry paid an annual percentage rate of approximately 28%. Working through the math, the dishwasher was $320 if bought outright. Henry paid $20 down at the time so he was financing $300 ($320 - $20) and made $340 (10 * $34) in payments over 10 months. To determine the rate, we use the formula for computing an annuity: PRINCIPAL = (PMT / PERIODICRATE) * (1 - (1 / ((1 + PERIODICRATE) ^ Periods))) Where: PRINCIPAL = amount borrowed = $300 PMT = periodic payment = $34 PERIODS = number of payments = 10 PERIODICRATE= the APR we are looking for divided by 12 Solving for PeriodicRate we get about 2.34%. Multiply by 12 and we get 28.09%. Using Microsoft Excel, we could use the "rate()" formula as follows: =rate(10,-34,300) = 2.34% * 12 = 28.09%

OpenStudy (mathstudent55):

@Catlover5925 Your answer is way off. How much total interest did he pay?

OpenStudy (mathstudent55):

He paid 10 * $34 = $340 He gave a down payment of $20, so he only financed $300. He paid $340 - $300 = $40 in total interest. As simple interest, this would be 40/300 * 100 = 13.33% in 10 months. In 12 months, this is 13.33% / 10 * 12 = 16% This is a simplified calculation using simple interest, but the closes choice by far is D.

OpenStudy (mathstudent55):

^closest

OpenStudy (anonymous):

Answer D is wrong.